# Stochastic Online Metric Matching: Adversarial is no Harder than Stochastic

### v1

@@ 2 Preliminaries
In our model, $n$ requests and $m$ servers are located in a metric space $(X,\delta)$, where $n\leq m$ and $\delta:X\times X\to{\mathbb{R}}_{\geq 0}$ is a cost function satisfying the triangle inequality. Denote the locations of servers as $S=\{s_{1},\ldots,s_{m}\}$ and the locations of requests as $R=\{r_{1},\ldots,r_{n}\}$. The locations of servers are adversarially chosen and all available at the beginning. The requests arrive one by one, and their locations are independently drawn from a known distribution $\mathbb{D}$ over $X$. Upon the arrival of a new request, we need to match it to an available server immediately and irrevocably, and then both the request and the matched server leave the system. The cost of matching a server $s$ and a request $r$ is $\delta(s,r)$, and we aim to minimize the total matching cost.
@@
For all algorithm ${\mathcal{A}}$ and set of servers $S$ with $|S|\geq n$, denote $\mathrm{cost}_{{\mathcal{A}}}(S,n)$ as the expected cost of ${\mathcal{A}}$ when the initial set of servers is $S$ and $n$ requests are randomly drawn from $\mathbb{D}$, and denote $\mathrm{cost}_{{\mathcal{A}}}(n)$ as the expected cost of ${\mathcal{A}}$ when $|S|=|R|=n$ and all servers and requests are randomly drawn from $\mathbb{D}$. Given a matching $M$ between two sets $S$ and $R$, for each matched element $s\in S$ (resp. $r\in R$), we use $M(s)$ (resp. $M(r)$) to denote the element in $R$ (resp. $S$) that $s$ (resp. $r$) is matched to.
@@
We define the smoothness property of distributions as follows, which requires distributions to admit an upper bounded density function.
@@
We say that a distribution $\mathbb{D}$ over $X$ is $\beta$-smooth if $\mathbb{D}$ admits a density function $\mu_{\mathbb{D}}$ such that $\mu_{\mathbb{D}}(x)\leq\beta$ for all $x\in X$.
@@ 2.1 Benchmark
We adopt the optimal offline solution, defined as the optimal matching cost without any future uncertainty, as our benchmark. For all set of servers $S$ and set of requests $R$ such that $|S|\geq|R|$, denote $\mathrm{OPT}(S,R)$ as the minimum cost of all matchings between $S$ and $R$ that match all requests. Given a set of servers $S$ with $|S|\geq n$ and a distribution for request arrival $\mathbb{D}$, the optimal offline solution on $S$ with $n$ requests randomly drawn from $\mathbb{D}$ is defined as $\mathrm{OPT}(S,n)=\mathbb{E}_{R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]$, where $R\sim\mathbb{D}^{n}$ denotes that the requests in $R$ are independently drawn from $\mathbb{D}$. Furthermore, define $\mathrm{OPT}(n)=\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]$.
We use competitive ratio and regret to measure the performance of online algorithms. For given metric space $(X,\delta)$ and distribution $\mathbb{D}$, we say that the competitive ratio of an algorithm ${\mathcal{A}}$ is $\alpha$ if for all concerned instances [^3] We require $|S|=n$ when considering balanced markets and $n\leq|S|\leq O(n)$ for unbalanced markets. $(S,n)$,
$\mathrm{cost}_{{\mathcal{A}}}(S,n)\leq\alpha\cdot\mathrm{OPT}(S,n)\,.$
Moreover, define the regret of an algorithm ${\mathcal{A}}$ for an instance $(S,n)$ as
@@
2.2 Optimal Offline Solutions for Euclidean Space
A considerable portion of this paper focuses on Euclidean space. In this subsection, we assume the metric space to be $[0,1]^{d}$ with the Euclidean distance and bound the optimal offline solutions for smooth distributions and general distributions. All the omitted proofs in this subsection can be found in Appendix A.
@@ 2.2 Optimal Offline Solutions for Euclidean Space
Next, we establish the lower bounds for smooth distributions in balanced markets.
@@
The lower bound in Lemma 3 for $d=1$ follows from the quantile mapping and the well-known $\Omega(\sqrt{n})$ lower bound for the optimal offline solution with uniform distribution, and the lower bounds for $d\geq 2$ rely on nearest-neighbor-distance.
Finally, we present the lower bounds for smooth distributions in unbalanced markets, which are proved by employing nearest-neighbor-distance as well.
@@
2.3 The $\mathsf{Simulate\text{-}Optimize\text{-}Assign\text{-}Repeat}$ Algorithm
The $\mathsf{Simulate\text{-}Optimize\text{-}Assign\text{-}Repeat}$ ($\mathsf{SOAR}$) algorithm is proposed by [CKKZ23] under the setting in which both servers and requests are stochastic. The $\mathsf{SOAR}$ algorithm, whose guarantee is stated below, will serve as an important building block of our algorithms and be invoked in a black-box manner.
@@ 2.3 The $\mathsf{Simulate\text{-}Optimize\text{-}Assign\text{-}Repeat}$ Algorithm
Suppose that $m=n$ and $\mathbb{D}$ is a distribution on $X$, where the cost function $\delta$ is not required to satisfy the triangle inequality. When servers and requests are randomly drawn from $\mathbb{D}$, an algorithm exists with an expected cost of $\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)$.
3 Reducing Adversarial to Stochastic
In this section, we present an algorithmic framework to transform an algorithm for stochastic servers and requests into an algorithm for adversarial servers and stochastic requests at a moderate cost.
@@ 3 Reducing Adversarial to Stochastic
Given a distribution $\mathbb{D}$ on a metric space $(X,\delta)$, assume the existence of an algorithm ${\mathcal{A}}$ such that $\mathrm{cost}_{{\mathcal{A}}}(n)\leq\alpha\cdot\mathrm{OPT}(n)$. Then, an algorithm ${\mathcal{A}}^{\prime}$ exists such that $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $|S|\geq n$. If we further have $|S|=n$, then $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq(2\alpha+1)\cdot\mathrm{OPT}(S,n)$.
3.1 Proof of Theorem 1
Assuming that requests are stochastic, we start by lower bounding the optimal offline solution in balanced markets with adversarial servers. In particular, we show that it is at least the optimal offline solution in balanced markets with stochastic servers and requests up to a constant factor.
@@ 3.1 Proof of Theorem 1
For all metric space $(X,\delta)$, for all distribution $\mathbb{D}$ over $X$ and set of servers $S^{\prime}$ with $|S^{\prime}|=n$, $\mathrm{OPT}(S^{\prime},n)\geq\frac{1}{2}\cdot\mathrm{OPT}(n)$.
@@
Given a set of servers $S^{\prime}$ with $|S^{\prime}|=n$, we claim that $\mathrm{OPT}(S,R)\leq\mathrm{OPT}(S,S^{\prime})+\mathrm{OPT}(S^{\prime},R)$ for all set of servers $S$ and set of requests $R$ with $|S|=|R|=n$. To see this, let $M_{1}$ be the perfect matching between $S$ and $S^{\prime}$ with the minimum cost and $M_{2}$ be the perfect matching between $S^{\prime}$ and $R$ with the minimum cost. Note that
$\mathrm{OPT}(S,S^{\prime})+\mathrm{OPT}(S^{\prime},R)$ $=\sum_{s^{\prime}\in S^{\prime}}\delta(M_{1}(s^{\prime}),s^{\prime})+\sum_{s^{\prime}\in S^{\prime}}\delta(s^{\prime},M_{2}(s^{\prime}))$
$\geq\sum_{s^{\prime}\in S^{\prime}}\delta(M_{1}(s^{\prime}),M_{2}(s^{\prime}))$
$\geq\mathrm{OPT}(S,R)\,,$
where the first inequality holds by the triangle inequality. Taking expectations on both sides over $S,R\sim\mathbb{D}^{n}$, we obtain
$\mathrm{OPT}(n)\leq\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,S^{\prime})+\mathrm{OPT}(S^{\prime},R)\right]=2\cdot\mathrm{OPT}(S^{\prime},n)\,,$
@@
Now, we are ready to prove Theorem 1. Given a distribution $\mathbb{D}$ over a metric space $(X,\delta)$, let ${\mathcal{A}}$ be an algorithm such that $\mathrm{cost}_{{\mathcal{A}}}(n)\leq\alpha\cdot\mathrm{OPT}(n)$, and we construct an algorithm ${\mathcal{A}}^{\prime}$ for adversarial servers and stochastic requests as follows. Given a set of servers $S$, we randomly generate a set of points $I\sim\mathbb{D}^{n}$, and let $M^{S,I}$ be the matching between $S$ and $I$ with the minimum cost that matches all points in $I$. Then, we run ${\mathcal{A}}$ with the set of initial servers being $I$. For each new request $r$, suppose that ${\mathcal{A}}$ matches it to $i\in I$, and we match $r$ to $M^{S,I}(i)\in S$ in ${\mathcal{A}}^{\prime}$.
Next, we analyze the expected cost of ${\mathcal{A}}^{\prime}$. By the triangle inequality, the expected cost of ${\mathcal{A}}^{\prime}$ is upper bounded by
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathbb{E}_{I\sim\mathbb{D}^{n}}\left[\sum_{i\in I}\delta(M^{S,I}(i),i)+\mathrm{cost}_{{\mathcal{A}}}(I,n)\right]=\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)\,,$
@@
4 Euclidean Space with Smooth Distributions

### v2

@@ 2 Model Setup
In our model, $n$ requests and $m$ servers are located in a metric space $(X,\delta)$, where $n\leq m$ and $\delta:X\times X\to{\mathbb{R}}_{\geq 0}$ is a cost function satisfying the triangle inequality. Denote the locations of servers as $S=\{s_{1},\ldots,s_{m}\}$ and the locations of requests as $R=\{r_{1},\ldots,r_{n}\}$. The servers are all available at the beginning with their locations known, and the requests arrive sequentially with their locations revealed only upon arrivals. When a new request arrives, we need to match it to an available server immediately and irrevocably, and then the matched server becomes unavailable. The cost of matching a server $s$ and a request $r$ is $\delta(s,r)$, and we aim to minimize the total matching cost.
@@
Depending on the context, we will assume the locations of servers and requests are either adversarially chosen or independently drawn from a known distribution $\mathbb{D}$ over $X$. For all algorithm ${\mathcal{A}}$, set of servers $S$ with $|S|\geq n$, and set of requests $R$ with $|R|=n$, denote $\mathrm{cost}_{{\mathcal{A}}}(S,R)$ as the cost of ${\mathcal{A}}$ with the initial set of servers being $S$ and the set of requests being $R$; denote $\mathrm{cost}_{{\mathcal{A}}}(S,n)$ as the expected cost of ${\mathcal{A}}$ when the initial set of servers is $S$ and $n$ requests are randomly drawn from $\mathbb{D}$; denote $\mathrm{cost}_{{\mathcal{A}}}(n)$ as the expected cost of ${\mathcal{A}}$ when $|S|=|R|=n$ and all servers and requests are randomly drawn from $\mathbb{D}$.
@@
Given a matching $M$ between two sets $S$ and $R$, for each matched element $s\in S$ (resp. $r\in R$), we use $M(s)$ (resp. $M(r)$) to denote the element in $R$ (resp. $S$) that $s$ (resp. $r$) is matched to. We define the smoothness property of distributions as follows, which requires distributions to admit an upper bounded density function.
@@
We say that a distribution $\mathbb{D}$ over $X$ is $\beta$-smooth if $\mathbb{D}$ admits a density function $\mu_{\mathbb{D}}$ such that $\mu_{\mathbb{D}}(x)\leq\beta$ for every $x\in X$.
@@ 2.1 Benchmark
We adopt the optimal offline cost, defined as the optimal matching cost without any future uncertainty, as our benchmark. For all set of servers $S$ and set of requests $R$ such that $|S|\geq|R|$, denote $\mathrm{OPT}(S,R)$ as the minimum cost of all matchings between $S$ and $R$ that match all requests. Given a set of servers $S$ with $|S|\geq n$ and a distribution $\mathbb{D}$ over $X$, the optimal offline cost on $S$ with $n$ requests randomly drawn from $\mathbb{D}$ is defined as $\mathrm{OPT}(S,n)=\mathbb{E}_{R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]$, where $R\sim\mathbb{D}^{n}$ denotes that the requests in $R$ are independently drawn from $\mathbb{D}$. Furthermore, define $\mathrm{OPT}(n)=\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]$.
We use competitive ratio and regret to measure the performance of online algorithms. When servers and requests are adversarial, for given metric space $(X,\delta)$, we say that the competitive ratio of an algorithm ${\mathcal{A}}$ is $\alpha$ if for all concerned instances [^3] We require $|S|=n$ when considering balanced markets and $n\leq|S|\leq O(n)$ for unbalanced markets. $(S,R)$, $\mathrm{cost}_{{\mathcal{A}}}(S,R)\leq\alpha\cdot\mathrm{OPT}(S,R)$. Similarly, when servers are adversarial and requests are stochastic, for given metric space $(X,\delta)$ and distribution $\mathbb{D}$, we say that the competitive ratio of an algorithm ${\mathcal{A}}$ is $\alpha$ if for all concerned instances $(S,n)$, $\mathrm{cost}_{{\mathcal{A}}}(S,n)\leq\alpha\cdot\mathrm{OPT}(S,n)$; moreover, define the regret of an algorithm ${\mathcal{A}}$ for a concerned instance $(S,n)$ as
@@
2.2 Optimal Offline Costs for Euclidean Space
A considerable portion of this paper focuses on Euclidean space. In this subsection, we assume the metric space to be $[0,1]^{d}$ with the Euclidean distance and bound the optimal offline costs for smooth distributions and general distributions. All the omitted proofs in this subsection can be found in Appendix A.
@@ 2.2 Optimal Offline Costs for Euclidean Space
Next, we establish lower bounds for smooth distributions in balanced markets.
@@
The lower bound in Lemma 3 for $d=1$ follows from the quantile mapping and the well-known $\Omega(\sqrt{n})$ lower bound for the optimal offline cost with uniform distribution, and the lower bounds for $d\geq 2$ rely on nearest-neighbor-distance.
Finally, we present lower bounds for smooth distributions in unbalanced markets, which are proved by employing nearest-neighbor-distance as well.
@@
2.3 The Simulate-Optimize-Assign-Repeat Algorithm
The Simulate-Optimize-Assign-Repeat (SOAR) algorithm is proposed by [CKKZ23] under the setting in which both servers and requests are stochastic. The SOAR algorithm, whose guarantee is stated below, will serve as an important building block of our algorithms for stochastic online metric matching and be invoked in a black-box manner.
@@ 2.3 The Simulate-Optimize-Assign-Repeat Algorithm
Suppose that $m=n$ and $\mathbb{D}$ is a distribution on $X$, where the cost function $\delta$ is not required to satisfy the triangle inequality. When all servers and requests are independently drawn from $\mathbb{D}$, an algorithm exists with an expected cost of $\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)$.
3 A Generic Approach for Beyond the Worst-Case Analysis
In this section, we present a generic algorithmic approach for online metric matching to go beyond the worst-case analysis. Then, we will demonstrate how to apply this approach to leverage the provided distributional information or predicted request locations.
@@ 3 A Generic Approach for Beyond the Worst-Case Analysis
Fix a metric space $(X,\delta)$ and an algorithm ${\mathcal{A}}$ for online metric matching. Then, there exists an algorithm ${\mathcal{A}}^{\prime}$ such that for all set of points $I$ given before the arrival of any request, set of servers $S$, and set of requests $R$ with $|S|\geq|I|=|R|$, $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)\leq\mathrm{OPT}(S,I)+\mathrm{cost}_{{\mathcal{A}}}(I,R)$. Moreover, if ${\mathcal{A}}$ is deterministic, then ${\mathcal{A}}^{\prime}$ is also deterministic.
Proof.
We construct the algorithm ${\mathcal{A}}^{\prime}$ as follows. Fix three sets of points $S,I,R$ as required. Let $M^{S,I}$ be the matching between $S$ and $I$ with the minimum cost that matches all points in $I$. Then, we run ${\mathcal{A}}$ with the set of initial servers being $I$. For each new request $r$, suppose that ${\mathcal{A}}$ matches it to $i\in I$, and ${\mathcal{A}}^{\prime}$ will match $r$ to $M^{S,I}(i)\in S$.
Since ${\mathcal{A}}^{\prime}$ does not use any randomness except for the invocation of ${\mathcal{A}}$, ${\mathcal{A}}^{\prime}$ is deterministic provided that ${\mathcal{A}}$ is deterministic. By the triangle inequality, the cost incurred by ${\mathcal{A}}^{\prime}$ is upper bounded by
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)\leq\sum_{i\in I}\delta(M^{S,I}(i),i)+\mathrm{cost}_{{\mathcal{A}}}(I,R)=\mathrm{OPT}(S,I)+\mathrm{cost}_{{\mathcal{A}}}(I,R)\,,$
concluding the proof. ∎
3.1 Stochastic Setting
We first apply Theorem 1 to the case where all request locations are independently drawn from a known distribution $\mathbb{D}$. We establish in the following theorem that an algorithm for stochastic servers and requests can be transformed into an algorithm for adversarial servers and stochastic requests in a black-box manner. In particular, for balanced markets, this transformation preserves the competitive ratio of the algorithm up to a constant factor.
Theorem 2.
Given a metric space $(X,\delta)$ and a distribution $\mathbb{D}$ over $X$, assume the existence of an algorithm ${\mathcal{A}}$ such that $\mathrm{cost}_{{\mathcal{A}}}(n)\leq\alpha\cdot\mathrm{OPT}(n)$. Then, an algorithm ${\mathcal{A}}^{\prime}$ exists such that $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $|S|\geq n$. If we further have $|S|=n$, then $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq(2\alpha+1)\cdot\mathrm{OPT}(S,n)$.
The rest of this subsection is devoted to proving Theorem 2. We start with proving the triangle inequality for $\mathrm{OPT}(\cdot,\cdot)$.
@@ 3.1 Stochastic Setting
For all sets of points $S,I,R$ with $|S|=|I|=|R|$, $\mathrm{OPT}(S,R)\leq\mathrm{OPT}(S,I)+\mathrm{OPT}(I,R)$.
@@
Given $S,I,R$ with $|S|=|I|=|R|$, let $M_{1}$ be the min-cost perfect matching between $S$ and $I$, and let $M_{2}$ be the min-cost perfect matching between $I$ and $R$. Note that
$\mathrm{OPT}(S,I)+\mathrm{OPT}(I,R)=\sum_{i\in I}\delta(M_{1}(i),i)+\sum_{i\in I}\delta(i,M_{2}(i))\geq\sum_{i\in I}\delta(M_{1}(i),M_{2}(i))\geq\mathrm{OPT}(S,R)\,,$
where the first inequality holds by the triangle inequality. ∎
Assuming that requests are stochastic, we start by lower bounding the optimal offline cost in balanced markets with adversarial servers. In particular, we show that it is at least the optimal offline cost in balanced markets with stochastic servers and requests up to a constant factor.
Lemma 7.
For all metric space $(X,\delta)$, distribution $\mathbb{D}$ over $X$, and set of servers $S$ with $|S|=n$, $\mathrm{OPT}(S,n)\geq\frac{1}{2}\cdot\mathrm{OPT}(n)$.
Proof.
Given a set of servers $S$ with $|S|=n$, by Lemma 6, we know that $\mathrm{OPT}(I,R)\leq\mathrm{OPT}(I,S)+\mathrm{OPT}(S,R)$ for all sets of points $I,R$ with $|I|=|R|=n$. Taking expectations on both sides over $I,R\sim\mathbb{D}^{n}$, we obtain
$\mathrm{OPT}(n)\leq\mathbb{E}_{I,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(I,S)+\mathrm{OPT}(S,R)\right]=2\cdot\mathrm{OPT}(S,n)\,,$
@@
Now, we are ready to prove Theorem 2. Given a metric space $(X,\delta)$ and a distribution $\mathbb{D}$ over $X$, let ${\mathcal{A}}$ be an algorithm such that $\mathrm{cost}_{{\mathcal{A}}}(n)\leq\alpha\cdot\mathrm{OPT}(n)$, and we construct an algorithm ${\mathcal{A}}^{\prime}$ for adversarial servers and stochastic requests as follows. Given as input a set of servers $S$ and a set of requests $R$ with $|S|\geq|R|=n$, we randomly generate a set of points $I\sim\mathbb{D}^{n}$ before the arrival of any request, and let ${\mathcal{A}}^{\prime}$ be the algorithm as per Theorem 2, which follows that
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)\leq\mathrm{OPT}(S,I)+\mathrm{cost}_{{\mathcal{A}}}(I,R)\,.$
Since $I$ and $R$ are independent, the expected cost of ${\mathcal{A}}^{\prime}$ is upper bounded by
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathbb{E}_{I,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,I)+\mathrm{cost}_{{\mathcal{A}}}(I,R)\right]=\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)\,,$
@@
3.2 Prediction Setting
In this subsection, we focus on balanced markets, i.e., $m=n$, and assume that the locations of servers and requests are adversarially chosen. In addition, we assume that algorithms are equipped with a prediction of request locations. We then apply Theorem 1 to give an algorithm whose performance smoothly depends on the prediction error, i.e., a consistent algorithm. More formally, the prediction is given in the form of a set of predicted request locations $R^{\prime}$ with $|R^{\prime}|=n$, which is provided before the arrival of any request, and the prediction error is defined as $\eta\triangleq\mathrm{OPT}(R^{\prime},R)$. We reiterate that the prediction error is not known by algorithms. Throughout, we slightly abuse notations and use $\mathrm{OPT}\triangleq\mathrm{OPT}(S,R)$ to denote the optimal offline cost.
Theorem 3.
Given a metric space $(X,\delta)$, assume that there exists an $\alpha$-competitive algorithm ${\mathcal{A}}$ that does not use predictions. Then, there exists an algorithm ${\mathcal{A}}^{\prime}$ such that, when equipped with a prediction of request locations with error $\eta$, ${\mathcal{A}}^{\prime}$ is $(1+\frac{\eta(\alpha+1)}{\mathrm{OPT}})$-competitive. Moreover, if ${\mathcal{A}}$ is deterministic, then ${\mathcal{A}}^{\prime}$ is also deterministic.
Proof.
We construct ${\mathcal{A}}^{\prime}$ as follows. Given as input a set of servers $S$, a set of requests $R$, and a prediction of request locations $R^{\prime}$ with $|S|=|R|=|R^{\prime}|$ and $\eta=\mathrm{OPT}(R,R^{\prime})$, let ${\mathcal{A}}^{\prime}$ be the algorithm as per Theorem 2 by setting $I=R^{\prime}$. It gives that,
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)$ $\leq\mathrm{OPT}(S,R^{\prime})+\mathrm{cost}_{{\mathcal{A}}}(R^{\prime},R)\leq\mathrm{OPT}(S,R)+\mathrm{OPT}(R,R^{\prime})+\mathrm{cost}_{{\mathcal{A}}}(R^{\prime},R)$
$\leq\mathrm{OPT}(S,R)+\mathrm{OPT}(R,R^{\prime})+\alpha\cdot\mathrm{OPT}(R^{\prime},R)=\mathrm{OPT}(S,R)+\eta(\alpha+1)\,,$
where the first inequality holds by the guarantee of Theorem 2, the second inequality holds by Lemma 6, and the third inequality holds due to the competitive ratio of ${\mathcal{A}}$. Finally, if ${\mathcal{A}}$ is deterministic, then ${\mathcal{A}}^{\prime}$ being deterministic follows from the guarantee of Theorem 2. ∎
4 Stochastic Setting: Euclidean Space with Smooth Distributions